# Global analysis of charm mixing parameters and determination of the CKM angle $\mathbf{\gamma}$

### v1

@@ Appendix D Fit results
$\phi_{2}^{M}[\degree]$ $2.4$$\pm 1.6$ $[-0.7,5.7]$ ϕ2Γ[°]superscriptsubscriptitalic-$\phi_{2}^{\Gamma}[\degree]$ $2.6$$\pm 1.2$ $[0.2,5.0]$
$F_{+}^{4\pi}$ $0.745$$\pm 0.010$ $[0.725,0.764]$ $x_{12}\simeq x[\%]$ $0.402$$\pm 0.034$ $[0.335,0.470]$
$y_{12}\simeq y[\%]$ $0.629$$\pm 0.021$ $[0.587,0.672]$ $F_{+}^{\pi\pi\pi^{0}}$ $0.973$$\pm 0.015$ $[0.944,1.000]$
$a_{D}^{KK}[\%]$ $0.085$$\pm 0.063$ $[-0.041,0.211]$ $a_{D}^{\pi\pi}[\%]$ $0.301$$\pm 0.070$ $[0.161,0.443]$
$F_{+}^{KK\pi^{0}}$ $0.738$$\pm 0.055$ $[0.629,0.847]$ $r_{D}^{K\pi}[\%]$ $5.845$$\pm 0.017$ $[5.812,5.878]$
$\delta_{D}^{K\pi}[\degree]$ $191.1$$\pm 2.9$ $[185.3,196.8]$ $F_{+}^{KK\pi\pi}$ $0.830$$\pm 0.071$ $[0.691,0.969]$
$r_{D}^{K3\pi}[\%]$ $5.564$$\pm 0.058$ $[5.448,5.682]$ $\delta_{D}^{K3\pi}[\degree]$ $157$$\pm 15$ $[127,187]$
$\kappa_{D}^{K3\pi}$ $0.478$$\pm 0.064$ $[0.346,0.606]$ $r_{D}^{K\pi\pi^{0}}[\%]$ $4.43$$\pm 0.11$ $[4.22,4.64]$
$\delta_{D}^{K\pi\pi^{0}}[\degree]$ $203.3$$\pm 9.0$ $[185.1,221.2]$ $\kappa_{D}^{K\pi\pi^{0}}$ $0.794$$\pm 0.040$ $[0.714,0.873]$
$r_{D}^{K^{0}_{S}K\pi}$ $0.6196$$\pm 0.0053$ $[0.6090,0.6304]$ $\delta_{D}^{K^{0}_{S}K\pi}[\degree]$ $13$$\pm 14$ $[-16,40]$
$\kappa_{D}^{K^{0}_{S}K\pi}$ $0.821$$\pm 0.094$ $[0.647,0.996]$ $|q/p|-1[\%]$ $-0.1$$\pm 1.6$ $[-3.3,3.1]$
ϕ2[°]subscriptitalic-$\phi_{2}[\degree]$ $-2.54$$\pm 0.94$ $[-4.44,-0.66]$ ϕ12[°]subscriptitalic-$\phi_{12}[\degree]$ $-0.2$$\pm 2.0$ $[-4.1,4.0]$
Table 10: Results for the charm part of the combination. For each of the parameters, we have reported the central value (“Value”) and the smallest intervals containing the $68.3\%$ (“Unc.”) and $95.4\%$ probabilities.
@@
$r_{B}^{DK}[\%]$ $9.99$$\pm 0.23$ $[9.53,10.46]$ $r_{B}^{D\pi}[10^{-3}]$ $5.05$$\pm 0.72$ $[3.78,6.79]$
$\mathbf{B^{\pm}}$ $\delta_{B}^{DK}[\degree]$ $128.7$$\pm 2.9$ $[122.6,134.2]$ $\delta_{B}^{D\pi}[\degree]$ $297$$\pm 12$ $[270,319]$
Parameters $r_{B}^{D^{*}K}$ $0.109$$\pm 0.011$ $[0.087,0.129]$ $r_{B}^{D^{*}\pi}[\%]$ $0.59$$\pm 0.34$ $[0,1.3]$
$\delta_{B}^{D^{*}K}[\degree]$ $-44.0$$\pm 6.7$ $[-58.6,-31.6]$ $\delta_{B}^{D^{*}\pi}[\degree]$ $51$$\pm 41$ $[-7,161]$
$r_{B}^{DK^{*\pm}}$ $0.098$$\pm 0.009$ $[0.078,0.115]$ $\delta_{B}^{DK^{*\pm}}[\degree]$ $49$$\pm 12$ $[27,79]$
$\kappa_{B}^{DK^{*\pm}}$ $0.957$$\pm 0.040$ $[0.856,1.000]$ $\frac{\mathcal{B}(B\to DK)}{\mathcal{B}(B\to D\pi)}[\%]$ $7.79$$\pm 0.37$ $[7.05,8.54]$
CKM Angles $\gamma[\degree]$ $68.4$$\pm 2.6$ $[63.2,73.5]$ $\beta[\degree]$ $22.33$$\pm 0.57$ $[21.19,23.48]$
$-2\beta_{s}[\degree]$ $-2.7$$\pm 1.1$ $[-4.9,-0.6]$
$\mathbf{B^{0}}$ $r_{B^{0}}^{DK^{*0}}$ $0.240$$\pm 0.012$ $[0.215,0.265]$ $\delta_{B^{0}}^{DK^{*0}}[\degree]$ $192.9$$\pm 4.9$ $[183.1,203.1]$
Parameters $\kappa_{B^{0}}^{DK^{*0}}$ $0.957$$\pm 0.023$ $[0.914,1.000]$ $r_{B^{0}}^{D\pi}$ $0.027$$\pm 0.012$ $[0.002,0.050]$
$\delta_{B^{0}}^{D\pi}[\degree]$ $25$$\pm 30$ $[-51,80]$
$r_{B^{0}_{s}}^{D_{s}K}$ $0.328$$\pm 0.035$ $[0.259,0.399]$ $\delta_{B^{0}_{s}}^{D_{s}K}[\degree]$ $-11.5$$\pm 6.0$ $[-23.7,0.6]$
$\mathbf{B^{0}_{s}}$ $r_{B^{0}_{s}}^{D_{s}K\pi\pi}$ $0.449$$\pm 0.078$ $[0.294,0.604]$ $\delta_{B^{0}_{s}}^{D_{s}K\pi\pi}[\degree]$ $-16$$\pm 13$ $[-41,10]$
Parameters $\kappa_{B^{0}_{s}}^{D_{s}K\pi\pi}$ $0.74$$\pm 0.14$ $[0.50,1.00]$ $r_{B^{0}_{s}}^{D\overline{K^{*0}}}$ $0.060$$\pm 0.030$ $[0,0.101]$
$\delta_{B^{0}_{s}}^{D\overline{K^{*0}}}[\degree]$ $80$$\pm 72$ $[-62,239]$ $\kappa_{B^{0}_{s}}^{D\overline{K^{*0}}}$ $0.3$$\pm 0.3$ $[0,0.9]$
@@
Table 11: Results for the beauty part of the combination. For each of the parameters, we have reported the central value (“Value”) and the smallest intervals containing the $68.3\%$ (“Unc.”) and $95.4\%$ probabilities.

### v3

@@ Appendix C Fit results
$\phi_{2}^{M}[\degree]$ $0.13$$\pm 0.70$ $[-1.27,1.54]$ ϕ2Γ[°]superscriptsubscriptitalic-$\phi_{2}^{\Gamma}[\degree]$ $2.1$$\pm 1.6$ $[-1.0,5.2]$
ϕ2[°]subscriptitalic-$\phi_{2}[\degree]$ $-1.5$$\pm 1.1$ $[-3.7,0.6]$ $|q/p|-1[\%]$ $-1.6$$\pm 1.5$ $[-4.5,1.4]$
$x_{12}\simeq x[\permil]$ $4.01$$\pm 0.43$ $[3.16,4.83]$ $y_{12}\simeq y[\permil]$ $6.10$$\pm 0.17$ $[5.77,6.45]$
$a_{D}^{KK}[\permil]$ $0.40$$\pm 0.53$ $[-0.63,1.43]$ $a_{D}^{\pi\pi}[\permil]$ $2.34$$\pm 0.60$ $[1.15,3.53]$
ϕ12[°]subscriptitalic-$\phi_{12}[\degree]$ $-2.0$$\pm 1.8$ $[-5.6,1.7]$ $\kappa_{D}^{K3\pi}$ $0.444$$\pm 0.092$ $[0.261,0.617]$
$r_{D}^{K3\pi}$ $0.05502$$\pm 0.00070$ $[0.05364,0.05636]$ $\delta_{D}^{K3\pi}[\degree]$ $164$$\pm 21$ $[126,208]$
$r_{D}^{K\pi}[\%]$ $5.8573$$\pm 0.0094$ $[5.8387,5.8758]$ $\delta_{D}^{K\pi}[\degree]$ $191.4$$\pm 2.4$ $[186.6,196.0]$
$\kappa_{D}^{K\pi\pi^{0}}$ $0.791$$\pm 0.039$ $[0.713,0.868]$ $r_{D}^{K\pi\pi^{0}}$ $0.04433$$\pm 0.00099$ $[0.04239,0.04626]$
$\delta_{D}^{K\pi\pi^{0}}[\degree]$ $201.0$$\pm 7.9$ $[185.2,216.3]$ $\kappa_{D}^{K^{0}_{S}K\pi}$ $0.824$$\pm 0.094$ $[0.652,0.995]$
$r_{D}^{K^{0}_{S}K\pi}$ $0.6192$$\pm 0.0052$ $[0.6090,0.6297]$ $\delta_{D}^{K^{0}_{S}K\pi}[\degree]$ $13$$\pm 14$ $[-15,39]$
$F^{4\pi}_{+}$ $0.7444$$\pm 0.0097$ $[0.7252,0.7635]$ $F^{KK\pi\pi}_{+}$ $0.756$$\pm 0.013$ $[0.732,0.781]$
$F^{\pi\pi\pi^{0}}_{+}$ $0.9417$$\pm 0.0040$ $[0.9337,0.9496]$ $F^{KK\pi^{0}}_{+}$ $0.641$$\pm 0.017$ $[0.608,0.674]$
$(\Delta Y^{KK}-\Delta Y^{\pi\pi})[\permil]$ $0.24$$\pm 0.26$ $[-0.27,0.75]$
Table 8: Results for the charm part of the combination. For each of the parameters, we have reported the central value (“Value”) and the smallest intervals containing at least $68\%$ (“Unc.”) and $95\%$ probabilities.
@@
CKM Angles $\gamma[\degree]$ $65.7$$\pm 2.5$ $[60.8,70.4]$ $\phi_{d}/2[\degree]$ $22.13$$\pm 0.60$ $[20.96,23.32]$
$\phi_{s}[\degree]$ $-3.36$$\pm 0.80$ $[-4.93,-1.79]$
$\mathbf{B^{\pm}}$ $\kappa_{B}^{DK^{*\pm}}$ $0.950$$\pm 0.043$ $[0.852,1.001]$ $r_{B}^{DK^{*\pm}}$ $0.0994$$\pm 0.0090$ $[0.0806,0.1162]$
Parameters $\delta_{B}^{DK^{*\pm}}[\degree]$ $51$$\pm 12$ $[30,80]$ $r_{B}^{DK}[\%]$ $9.80$$\pm 0.20$ $[9.42,10.19]$
$\delta_{B}^{DK}[\degree]$ $127.0$$\pm 2.5$ $[121.9,131.8]$ $r_{B}^{D\pi}[\permil]$ $5.18$$\pm 0.56$ $[4.17,6.38]$
$\delta_{B}^{D\pi}[\degree]$ $299.8$$\pm 8.5$ $[282.0,315.1]$ $r_{B}^{D^{*}K}$ $0.1078$$\pm 0.0088$ $[0.0905,0.1251]$
$\delta_{B}^{D^{*}K}[\degree]$ $-47.6$$\pm 5.9$ $[-60.0,-36.7]$ $r_{B}^{D^{*}\pi}$ $0.0058$$\pm 0.0030$ $[0.0001,0.0123]$
$\delta_{B}^{D^{*}\pi}[\degree]$ $39$$\pm 31$ $[-4,133]$ $\frac{\mathcal{B}(B\to DK)}{\mathcal{B}(B\to D\pi)}$ $0.0782$$\pm 0.0038$ $[0.0707,0.0857]$
$r_{B}^{DK\pi\pi}$ $0.048$$\pm 0.031$ $[0.000,0.096]$ $\delta_{B}^{DK\pi\pi}[\degree]$ $[-226,-154]\cup[-75,69]~{}@~{}68.5\%$ prob.
$\kappa_{B}^{DK\pi\pi}$ $0.28$$\pm 0.28$ $[0.00,0.92]$ $r_{B}^{D\pi\pi\pi}$ $0.0099$ ±$\pm 0.0099$ $[0.0000,0.0365]$
$\delta_{B}^{D\pi\pi\pi}[\degree]$ $[-204,-100]\cup[-72,11]$${}68.9\%$ prob. $\kappa_{B}^{D\pi\pi\pi}$ $0.23$$\pm 0.23$ $[0.00,0.88]$
$\mathbf{B^{0}}$ $r_{B^{0}}^{D\pi}$ $0.0229$$\pm 0.0090$ $[0.0058,0.0439]$ $\delta_{B^{0}}^{D\pi}[\degree]$ $32$$\pm 31$ $[-43,75]$
Parameters $\kappa_{B^{0}}^{DK^{*0}}$ $0.933$$\pm 0.025$ $[0.885,0.982]$ $r_{B^{0}}^{DK^{*0}}$ $0.231$$\pm 0.015$ $[0.200,0.259]$
$\delta_{B^{0}}^{DK^{*0}}[\degree]$ $192.6$$\pm 6.0$ $[180.9,205.2]$ $r_{B^{0}}^{D^{*}\pi}$ $0.0244$$\pm 0.0083$ $[0.0099,0.0476]$
$\delta_{B^{0}}^{D^{*}\pi}[\degree]$ $-34$$\pm 31$ $[-74,41]$ $r_{B^{0}}^{D\rho}$ $0.060$ ±$\pm 0.060$ $[0.000,0.261]$
$\delta_{B^{0}}^{D\rho}[\degree]$ $-77$$\pm 27$ $[-265,93]$
$r_{B^{0}_{s}}^{D_{s}K}$ $0.331$$\pm 0.035$ $[0.264,0.402]$ $\delta_{B^{0}_{s}}^{D_{s}K}[\degree]$ $-11.0$$\pm 6.0$ $[-22.8,0.8]$
$\mathbf{B^{0}_{s}}$ $\kappa_{B^{0}_{s}}^{D_{s}K\pi\pi}$ $0.76$$\pm 0.14$ $[0.51,1.00]$ $r_{B^{0}_{s}}^{D_{s}K\pi\pi}$ $0.453$$\pm 0.077$ $[0.301,0.605]$
Parameters $\delta_{B^{0}_{s}}^{D_{s}K\pi\pi}[\degree]$ $-14$$\pm 12$ $[-39,10]$ $\kappa_{B^{0}_{s}}^{D\overline{K^{*0}}}$ $0.40$$\pm 0.30$ $[0.00,0.92]$
$r_{B^{0}_{s}}^{D\overline{K^{*0}}}$ $0.057$$\pm 0.029$ $[0.000,0.098]$ $\delta_{B^{0}_{s}}^{D\overline{K^{*0}}}[\degree]$ $72$$\pm 58$ $[-44,215]$
@@
Table 9: Results for the beauty part of the combination. For each of the parameters, we have reported the central value (“Value”) and the smallest intervals containing at least $68\%$ (“Unc.”) and $95\%$ probabilities. For the phases $\delta_{B}^{DK\pi\pi}$ and $\delta_{B}^{D\pi\pi\pi\pi}$, we report the two disjoint intervals at $68.5\%$ and $68.9\%$ probabilities, respectively.
Parameter Value Unc. Exp. Parameter Value Unc. Exp.
$(\left\langle\tau\right\rangle^{KK}_{\Delta A_{D}^{\mathrm{CP}}}+\left\langle\tau\right\rangle^{\pi\pi}_{\Delta A_{D}^{\mathrm{CP}}})/2$ $1.73$$\pm 0.10$ LHCb Run2, $\pi$-tag LHCb:2019hro $\left\langle\tau\right\rangle^{KK}_{\Delta A_{D}^{\mathrm{CP}}}-\left\langle\tau\right\rangle^{\pi\pi}_{\Delta A_{D}^{\mathrm{CP}}}$ $0.1350$$\pm 0.0020$ LHCb Run2, $\pi$-tag LHCb:2019hro
$(\left\langle\tau\right\rangle^{KK}_{\Delta A_{D}^{\mathrm{CP}}}+\left\langle\tau\right\rangle^{\pi\pi}_{\Delta A_{D}^{\mathrm{CP}}})/2$ $1.210$$\pm 0.010$ LHCb Run2, $\mu$-tag LHCb:2019hro $\left\langle\tau\right\rangle^{KK}_{\Delta A_{D}^{\mathrm{CP}}}-\left\langle\tau\right\rangle^{\pi\pi}_{\Delta A_{D}^{\mathrm{CP}}}$ $-0.0030$$\pm 0.0010$ LHCb Run2, $\mu$-tag LHCb:2019hro
$\left\langle t\right\rangle^{KK}_{A_{D}^{\mathrm{CP}}}[\mathrm{ps}]$ $0.7315$$\pm 0.0020$ LHCb Run2, $D^{+}$-cal LHCb:2022lry $\left\langle t\right\rangle^{KK}_{A_{D}^{\mathrm{CP}}}[\mathrm{ps}]$ $0.6868$$\pm 0.0014$ LHCb Run2, $D_{s}^{+}$-cal LHCb:2022lry
$\left\langle\tau\right\rangle^{KK}_{\Delta A_{D}^{\mathrm{CP}}}$ $2.152$$\pm 0.016$ LHCb Run1, $\pi$-tag LHCb:2016nxk $\left\langle\tau\right\rangle^{\pi\pi}_{\Delta A_{D}^{\mathrm{CP}}}$ $2.037$$\pm 0.015$ LHCb Run1, $\pi$-tag LHCb:2016nxk
$\left\langle\tau\right\rangle^{KK}_{\Delta A_{D}^{\mathrm{CP}}}$ $1.0820$$\pm 0.0041$ LHCb Run1, $\mu$-tag LHCb:2016nxk $\left\langle\tau\right\rangle^{\pi\pi}_{\Delta A_{D}^{\mathrm{CP}}}$ $1.0679$$\pm 0.0041$ LHCb Run1, $\mu$-tag LHCb:2016nxk
$\left\langle\tau\right\rangle^{KK}_{A_{D}^{\mathrm{CP}}}$ $2.239$$\pm 0.019$ LHCb Run1, $\pi$-tag LHCb:2016nxk $\left\langle\tau\right\rangle^{KK}_{A_{D}^{\mathrm{CP}}}$ $1.0510$$\pm 0.0041$ LHCb Run1, $\mu$-tag LHCb:2016nxk
$\left\langle\tau\right\rangle^{KK}_{A_{D}^{\mathrm{CP}}}$ $2.65$$\pm 0.03$ CDF CDF:2011ejf $\left\langle\tau\right\rangle^{\pi\pi}_{A_{D}^{\mathrm{CP}}}$ $2.40$$\pm 0.03$ CDF CDF:2011ejf
Table 10: Parameters corresponding to the average decay times entering the integrated CP asymmetries $A_{D}^{\mathrm{CP}}(f,t)$ in Eq. (29) and $\Delta A_{D}^{\mathrm{CP}}$ in Eq. (30). The column Exp. lists the name of the experiment, the tagging (-tag) mode employed, the calibration (-cal) method, and the reference describing the measurement.